1. A store manager wishes to find out whether there is a relationship between the age of her employees and the number of sick days they take each year. The data for the sample are shown. Age ()18 26 39 48 53 58 Days () 16 12 9 56 2 Draw the scatterplot and describe the association between the two variables. a) b) Is there a significant (nonzero) linear relationship between the age of the employees and the number of sick days? Use 5% significance level. [ldentify the null hypothesis, altemative hypothesis, test statistic, critical value from Table A-6, conclusion about the null hypothesis, and final conclusion. -45- MATH 160 Last Name First Name: Student ID # c) If there is a significant linear relationship, find the regression line. d) Find the best predicted number of sick days for an employee at 45 year of age.
Answer #1

Question 1

Here to calculate regression coefficient or the regression line coefficient, we require the value of \sum x , \sum y , \sum xy , \sum x^2 , \sum y^2

Age(x) Days(y) x^2 y^2 xy
18 16 324 256 288
26 12 676 144 312
39 9 1521 81 351
48 5 2304 25 240
53 6 2809 36 318
58 2 3364 4 116
Sum 242 50 10998 546 1625

Here as we can see the direction of scatter plot is negative and relationship seems strong.

(b) Here regression coefficient

r = [nΣxy - ΣxΣy]/ sqrt [n(Σx2 - (Σx)2] [n(Σy2 - (Σy)2]

r = [6 * 1625 - 242 * 50]/ sqrt [(6 * 10998 - 2422) * (6 * 546 - 502)]

r = -2350/ sqrt [7424 * 776]

r = -2350/2400.31 = -0.9791

Here for n = 6

dF = n -2 = 6 - 2 = 4

t = r sqrt [(n-2) / (1 -r2 )] = -0.9791 * sqrt [(6 - 2)/ (1 - 0.97912)] = -9.6283

Here critical value for dF = 4

tcritical = 2.7764

Here t > tcritical  so we will reject the null hypothesis and conclude that there is significant relationship between the variables.

(c)

y^ = a + bx

a = [(Σy) (Σx2 ) - (Σx) (Σxy)]/ [ n (Σx2 ) - (Σx)2 ]

a = [50 * 10998 - 242 * 1625]/ [6 * 10998 - 2422]

a = 21.1005

b =  [ n(Σxy) - (Σx)((Σy)]/ [ n (Σx2 ) - (Σx)2 ]

b = [6 * 1625 - 242 * 50]/ [6 * 10998 - 2422]

b= -2350/7424 = -0.3165

y^ = 21.1005 - 0.3165 x

(d) Here x = 45

y^ = 21.1005 - 0.3165 * 45 = 6.86 days